An airplane wing with an average cross sectional area of 10 m2 experiences a lift force of 5000 N. What is the difference in air pressure, on the average, between the bottom and top of the wing?




1
Expert's answer
2022-03-29T05:33:37-0400

The difference in air pressure

"P=\frac{F}{A}\\
=\rm\frac{5000\: N}{10\: m^2}=500\:Pa"